A river is of width 120m which flows at a speed of 8m/s. If a man swims with a speed of 5m/s at an angle 127° with the stream, his drift on reaching other bank is

Grade:12

2 Answers

Milind Rokade
6 years ago
Velocity in x direction is 8-3=5m/s             Velocity in y direction is 4m/s                                                      Time taken is 120/4=30 second  drift is 5*30=150m
sweta
4 years ago
velocity along x axis=5
velocity along y axis is 4
time taken=30seconds
drift=5*30=150
this is the answer 
hope it helps you
and make sure to mark a 5 star and an upvote